# Balanced Image Stylization with Style Matching Score

### v1

@@ Abstract
We present Style Matching Score (SMS), a novel optimization method for image stylization with diffusion models. Unlike existing efforts, our method reframes image stylization as a style distribution matching problem. The target style distribution is estimated from off-the-shelf style-dependent LoRAs via carefully designed score functions. To preserve content information adaptively, we propose Progressive Spectrum Regularization, which operates in the frequency domain to guide stylization progressively from low-frequency layouts to high-frequency details. In addition, we devise a Semantic-Aware Gradient Refinement technique that leverages relevance maps derived from diffusion semantic priors to selectively stylize semantically important regions. The proposed optimization formulation extends stylization from pixel space to parameter space, readily applicable to lightweight feedforward generators for efficient one-step stylization. SMS effectively balances style alignment and content preservation, outperforming state-of-the-art approaches, verified by extensive experiments.
@@ 1 Introduction
Image stylization transforms a content image to adopt a specific visual style, with applications in art design [5], virtual reality [10], and content creation [26, 4, 27]. Early exemplar-based methods like Neural Style Transfer (NST) [5, 14] capture style using statistical correlations in CNN feature maps of a style image but are limited to single-image styles and are heavily content-influenced. Collection-based methods using GANs [40, 2] model style as distributions over datasets for more realistic transfer but require large, well-curated datasets and face issues like mode collapse and training instability.
@@ 2 Related Work
Image Style Transfer aims to synthesize images in artistic styles and has significant practical value. GAN-based methods [40, 2, 13] enable realistic stylization by learning style distributions from datasets, which however requires extensive data and are challenging to train. Recent work leverages pretrained diffusion models (DMs) for style transfer, categorized as follows. 1) Prompts-Guide Stylization, represented by FreeStyle [6], offers flexibility but struggles with ambiguous or imprecise style control due to the abstract nature of text guidance. 2) Exemplar-Based Stylization typically manipulates self-attention or cross-attention layers to transfer style from a reference image. InST [39] maps style images to textual embeddings for conditions. StyleID [3] maintains content integrity through query preservation and AdaIN [12]. InstantStyle-Plus [31] injects style-specific features into intermediate model layers. Single-style exemplars, however, lead to over-stylization, obscuring semantic content. 3) Tuning-Based Stylization fine-tunes DMs on target style distributions, which produces high-quality results but is computationally costly. Style-dependent LoRA adapt models to specific styles with less data and capture nuanced style distributions. However, they typically use ControlNet [37] with basic edge maps from the source image for image-to-image stylization, which may lose details. To overcome these limitations, we propose a distribution-level style matching loss, fully leveraging diffusion priors for semantic-aware, identity-preserving and applicable stylization.
@@ 4 Feed-Forward Stylization with SMS
As an applications of SMS, we present a detailed data-free stylizaton pipeline using a lightweight feed-forward network $G_{\theta}$ (see Figure 4). Given a set of real images $\{x_{i}\}_{i=1}^{N}\in p_{\text{real}}$ and a target style domain $p_{\text{style}}$ modeled by a pretrained $\epsilon_{\text{style}}$, our goal is to learn a mapping $G_{\theta}:p_{\text{text}}\rightarrow p_{\text{style}}$ that transforms input images into the target style domain.
Reconstruction Warmup. To ensure that $G_{\theta}$ maintains content fidelity and accelerates convergence, we initialize $G_{\theta}$ by training it to reconstruct the input images. This warmup phase aligns the initial generated distribution $p_{G_{\theta}}$ with the real image distribution $p_{\text{real}}$, providing a solid foundation for subsequent training. The reconstruction loss is defined as: The reconstruction loss is defined as:
@@ 5.3 Ablation Studies
Timestep Sampling Strategy. We compare our adaptive narrowing sampling strategy with random and naive linear decreasing sampling under identical settings. Random sampling causes blurriness (see Figure 7(g)), while naive linear decreasing sampling fails to preserve local identity, such as a red scard-like artifact appearing around the dog’s neck (see Figure 7(h)). In contrast, our method produces sharp, detailed stylization while preserving content fidelity. The quantitative results in Table 4 clearly demonstrate the effectiveness of each proposed module. Additional ablation studies are provided in the appendix.
@@ 6 Conclusion
The document provides supplementary information not elaborated on in our main paper due to space constraints: implementation details (Section A), derivation of the Style Matching Objective (Section B), spectrum-based style analysis (Section C), additional optimization-based comparisons (Section D), further studies (Section E), more qualitative comparisons (Section F), additional results (Section G), limitations (Section H), and broader impact (Section I).
@@ 22      Sample $\epsilon\sim\mathcal{N}(0,\mathbf{I})$
A.2 Style Data
We leverage the off-the-shelf style LoRA from https://civitai.com/ to support diverse artistic styles. To ensure fairness when comparing with baseline methods that utilize different style representations, we carefully adopt the following adaptations: 1) Text-driven (e.g., FreeStyle [6], DDS [7]): Use descriptive text prompts to capture the style. 2) Exemplar-guided (e.g., StyleID [3], InstantStyle-Plus [31]): Source reference images from training data. 3) Collection-based (e.g., Style-LoRA): Use Exactly the same LoRA.
A.3 Training time
Table 5 reports the per-image runtime (seconds) for baselines on an L40 GPU under default settings. Although some baselines require only forward steps, they require additional processing steps such as DDIM [29] inversion and substantial preparation (e.g., ControlNet [37] training for InstantStyle-Plus [31] and Style-LoRA). In contrast, our SMS runs without any model-specific preparation, achieving a comparable overall runtimes. Furthermore, the optimization-based nature of SMS enables it to extend to more complex, parameterized representations, which is not straightforward with other methods.
Table 5: Per-Image Runtime Comparison.
@@ E.1 Identity Loss Variant Study
Figure 11: Effects of the loss weight $\lambda$.
@@ E.2 Effects of Loss Weight $\lambda$ Study
The strength of the explicit identity regularization term is determined by the loss weight $\lambda$. As shown in Figure 11, increasing $\lambda$ enhances content fidelity, while reducing it allows for stronger stylization. It provides a control to adjust the level of stylization.

### v2

@@ Abstract
We present Style Matching Score (SMS), a novel optimization method for image stylization with diffusion models. Balancing effective style transfer with content preservation is a long-standing challenge. Unlike existing efforts, our method reframes image stylization as a style distribution matching problem. The target style distribution is estimated from off-the-shelf style-dependent LoRAs via carefully designed score functions. footnotetext: Corresponding author. To preserve content information adaptively, we propose Progressive Spectrum Regularization, which operates in the frequency domain to guide stylization progressively from low-frequency layouts to high-frequency details. In addition, we devise a Semantic-Aware Gradient Refinement technique that leverages relevance maps derived from diffusion semantic priors to selectively stylize semantically important regions. The proposed optimization formulation extends stylization from pixel space to parameter space, readily applicable to lightweight feedforward generators for efficient one-step stylization. SMS effectively balances style alignment and content preservation, outperforming state-of-the-art approaches, verified by extensive experiments. Code: https://github.com/showlab/SMS.
@@ 1 Introduction
Image stylization transforms a content image to adopt a specific visual style, with applications in art design [6], virtual reality [11], and content creation [27, 5, 28, 32]. Early exemplar-based methods like Neural Style Transfer (NST) [6, 15] capture style using statistical correlations in CNN feature maps of a style image but are limited to single-image styles and are heavily content-influenced. Collection-based methods using GANs [42, 3] model style as distributions over datasets for more realistic transfer but require large, well-curated datasets and face issues like mode collapse and training instability.
@@ 2 Related Work
Image Style Transfer aims to synthesize images in artistic styles and has significant practical value. GAN-based methods [42, 3, 14] enable realistic stylization by learning style distributions from datasets, which however requires extensive data and are challenging to train. Recent work leverages pretrained diffusion models (DMs) for style transfer, categorized as follows. 1) Prompts-Guide Stylization, represented by FreeStyle [7], offers flexibility but struggles with ambiguous or imprecise style control due to the abstract nature of text guidance. 2) Exemplar-Based Stylization typically manipulates self-attention or cross-attention layers to transfer style from a reference image. InST [41] maps style images to textual embeddings for conditions. StyleID [4] maintains content integrity through query preservation and AdaIN [13]. InstantStyle-Plus [33] injects style-specific features into intermediate model layers. Single-style exemplars, however, lead to over-stylization, obscuring semantic content. 3) Tuning-Based Stylization fine-tunes DMs on target style distributions, which produces high-quality results but is computationally costly. Style-dependent LoRA adapt models to specific styles with less data and capture nuanced style distributions [1]. However, they typically use ControlNet [39] with basic edge maps from the source image for image-to-image stylization, which may lose details. To overcome these limitations, we propose a distribution-level style matching loss, fully leveraging diffusion priors for semantic-aware, identity-preserving and applicable stylization.
@@ 4 Feed-Forward Stylization with SMS
As an applications of SMS, we present a detailed data-free stylizaton pipeline using a lightweight feed-forward network $G_{\theta}$ (see Figure 4). Given a set of real images $\{x_{i}\}_{i=1}^{N}\in p_{\text{real}}$ and a target style domain $p_{\text{style}}$ modeled by a pretrained $\epsilon_{\text{style}}$, our goal is to learn a mapping $G_{\theta}:p_{\text{text}}\rightarrow p_{\text{style}}$ that transforms input images into the target style domain. Reconstruction Warmup. To ensure that $G_{\theta}$ maintains content fidelity and accelerates convergence, we initialize $G_{\theta}$ by training it to reconstruct the input images. This warmup phase aligns the initial generated distribution $p_{G_{\theta}}$ with the real image distribution $p_{\text{real}}$, providing a solid foundation for subsequent training. The reconstruction loss is defined as: The reconstruction loss is defined as:
@@ 5.3 Ablation Studies
Timestep Sampling Strategy. We compare our adaptive narrowing sampling strategy with random and naive linear decreasing sampling under identical settings. Random sampling causes blurriness (see Figure 7(f)), while naive linear decreasing sampling fails to preserve local identity, such as a red scard-like artifact appearing around the dog’s neck (see Figure 7(g)). In contrast, our method produces sharp, detailed stylization while preserving content fidelity. The quantitative results in Table 4 clearly demonstrate the effectiveness of each proposed module. Additional ablation studies are provided in the appendix.
@@ 6 Conclusion
Acknowledgment. This study is supported by the Ministry of Education, Singapore, under the Academic Research Fund Tier 1 (FY2023). Yuxin Jiang is supported by the A*STAR ACIS Scholarship.
@@
The document provides supplementary information not elaborated on in our main paper due to space constraints: implementation details (Section A), derivation of the Style Matching Objective (Section B), spectrum-based style analysis (Section C), additional optimization-based comparisons (Section D), further studies (Section E), more qualitative comparisons (Section F), additional results (Section G), limitations (Section H), and broader impact (Section I). Code: https://github.com/showlab/SMS.
@@ A.1 SMS Training Procedure
A.2 Style Data
We leverage the off-the-shelf style LoRA from Civitai [1] to support diverse artistic styles. To ensure fairness when comparing with baseline methods that utilize different style representations, we carefully adopt the following adaptations: 1) Text-driven (e.g., FreeStyle [7], DDS [8]): Use descriptive text prompts to capture the style. 2) Exemplar-guided (e.g., StyleID [4], InstantStyle-Plus [33]): Source reference images from training data. 3) Collection-based (e.g., Style-LoRA): Use Exactly the same LoRA.
A.3 Training time
Table 5 reports the per-image runtime (seconds) for baselines on an NVIDIA L40 GPU under default settings. Although some baselines require only forward steps, they require additional processing steps such as DDIM [30] inversion and substantial preparation (e.g., ControlNet [39] training for InstantStyle-Plus [33] and Style-LoRA). In contrast, our SMS runs without any model-specific preparation, achieving a comparable overall runtimes. Furthermore, the optimization-based nature of SMS enables it to extend to more complex, parameterized representations, which is not straightforward with other methods.
@@ 22      Sample $\epsilon\sim\mathcal{N}(0,\mathbf{I})$
Table 5: Image stylization per-image runtime comparison.
@@ E.1 Identity Loss Variant Study
Figure 11: Effects of the loss weight $\lambda$. The first row shows results in kids illustration style, and the last two rows show results in Ghibli style.
@@ E.2 Effects of Loss Weight $\lambda$ Study
The strength of the explicit identity regularization term is determined by the loss weight $\lambda$. As shown in Figure 11, increasing $\lambda$ enhances content fidelity, while reducing it allows for stronger stylization, demonstrating a clear trade-off between style and content. It provides a user-controllable knob for adjusting the stylization strength.